OpenStudy (anonymous):

I need A Bit of Help ^-^ (I'll Fan) If Anyone Could Thank You <3 Solve: 7x - 2(x + 1) = 6x + 14

OpenStudy (anonymous):

Simplifying 7x + -2(x + 1) = 6x + 14 Reorder the terms: 7x + -2(1 + x) = 6x + 14 7x + (1 * -2 + x * -2) = 6x + 14 7x + (-2 + -2x) = 6x + 14 Reorder the terms: -2 + 7x + -2x = 6x + 14 Combine like terms: 7x + -2x = 5x -2 + 5x = 6x + 14 Reorder the terms: -2 + 5x = 14 + 6x Solving -2 + 5x = 14 + 6x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-6x' to each side of the equation. -2 + 5x + -6x = 14 + 6x + -6x Combine like terms: 5x + -6x = -1x -2 + -1x = 14 + 6x + -6x Combine like terms: 6x + -6x = 0 -2 + -1x = 14 + 0 -2 + -1x = 14 Add '2' to each side of the equation. -2 + 2 + -1x = 14 + 2 Combine like terms: -2 + 2 = 0 0 + -1x = 14 + 2 -1x = 14 + 2 Combine like terms: 14 + 2 = 16 -1x = 16 Divide each side by '-1'. x = -16 Simplifying x = -16

OpenStudy (anonymous):

x = -16

OpenStudy (anonymous):

Thank you (: